On a normal day in the summer you can drive your 1050 kg car at a constant speed around

a corner whose radius of curvature is 12m without problems. The engineering staff at

CalTrans have determined that the maximum frictional force between your tires and the

road is 8100N under these circumstances.



What is the greatest speed that you can travel around the corner without

slipping?


1
Expert's answer
2021-10-05T10:05:57-0400

"F=mv^2/r\to v=\sqrt{Fr/m}=\sqrt{8100\cdot 12/1050}=9.62\ (m/s)" . Answer